The lateral area of a cone is 32pi and the slant height is 8. Find the radius, height, total area, and volume.
Answer by mananth(16946)


The lateral area of a cone is 32pi and the slant height is 8. Find the radius, height, total area, and volume.
lateral area = 32 pi
32 pi = pi r^2
r =sqrt(32) = 4 sqrt(2) ( radius)
Total surface area = pi *r^2 + pi*r*l l=mslant height)
TSA = pi*32*+pi*16sqrt(2)*8
TSA = 32pi(1+4sqrt(2))